### v2

@@ Abstract
The increasing demand for generative AI as Large Language Models (LLMs) services has driven the need for specialized hardware architectures that optimize computational efficiency and energy consumption. This paper evaluates the performance of the Tenstorrent Grayskull e75 RISC-V accelerator for basic linear algebra kernels at reduced numerical precision, a fundamental operation in LLM computations. We present a detailed characterization of Grayskull’s execution model, grid size, matrix dimensions, data formats, and numerical precision impact computational efficiency. Furthermore, we compare Grayskull’s performance against state-of-the-art architectures with tensor acceleration, including Intel Sapphire Rapids processors and two NVIDIA GPUs (V100 and A100). Whilst NVIDIA GPUs dominate raw performance, Grayskull demonstrates a competitive trade-off between power consumption and computational throughput, reaching a peak of 1.55 TFLOPs/Watt with BF16.
@@ 5.2 Offload and execution model
After the first execution, the program does not need to be re-compiled, leading to a significant speedup in subsequent runs for both the tiling and the MatMul operation run. With respect to the tiling in the first-run plot, the tiling of the first input (in0) is three orders of magnitude greater than the tiling of the second matrix (in1), as the tiling of the second matrix will reuse the already compiled tiling program for the first one. By calculating the compilation time as the difference between the execution times of the first and subsequent executions, we obtain that the tiling kernel requires 296 ms to compile, regardless of the matrix dimensions. Whilst, the tiling execution time ranges from 351 $\mu$s for the 256x256 matrix to 2256 $\mu$s for a 4096x4096 matrix. On the other hand, the MatMul kernel requires 620 ms for compilation, with executing times ranging from 328 $\mu$s to 4783 $\mu$s for the same matrix sizes.
@@ 5.3 Performance of the MatMul kernel
3(b) shows the speedup of the MatMul kernel for different sets of cores, ranging from 1 core to 64 cores, corresponding respectively to the grid core selection $(0,0)$ and $(8,8)$.
@@ 5.4.1 Performance comparison
The obtained performance results have been compared against the PyTorch MatMul kernel execution on three SoA architectures: two GPUs, the NVIDIA A100 SXM4 40GB (w. peak BF16 throughput of 312 TFLOPs) and the NVIDIA V100S PCIe 32GB (w. peak FP16 throughput of 32 TFLOPs), and the Intel Sapphire Rapids server processor (w. peak BF16 throughput of 229 TFLOPs) [^2] Calculated considering AMX’s 1024 FLOPS/Cycle, 112 cores and 2.0 GHz frequency. The comparison is shown in 5(a). The Sapphire Rapids Intel processor is equipped with the Intel Advanced Matrix Extensions (AMX) to optimize MatMul execution in BF16 and INT8 formats, which are leveraged by the PyTorch’s MatMul kernel. The NVIDIA GPUs obtain remarkable performance thanks to the introduction of Tensor Cores, which accelerate the MatMul computation using BF16 and FP16 formats.

### v3

@@ Abstract
The increasing demand for generative AI as Large Language Models (LLMs) services has driven the need for specialized hardware architectures that optimize computational efficiency and energy consumption. This paper evaluates the performance of the Tenstorrent Grayskull e75 RISC-V accelerator for basic linear algebra kernels at reduced numerical precision, a fundamental operation in LLM computations. We present a detailed characterization of Grayskull’s execution model, grid size, matrix dimensions, data formats, and numerical precision impact on computational efficiency. Furthermore, we compare Grayskull’s performance against state-of-the-art architectures with tensor acceleration, including Intel Sapphire Rapids processors and two NVIDIA GPUs (V100 and A100). Whilst NVIDIA GPUs dominate raw performance, Grayskull demonstrates a competitive trade-off between power consumption and computational throughput, reaching a peak of 1.55 TFLOPs/Watt with BF16.
@@ 5.2 Offload and execution model
After the first execution, the program does not need to be re-compiled, leading to a significant speedup in subsequent runs for both the tiling and the MatMul operation run. With respect to the tiling in the first-run plot, the tiling of the first input (in0) is three orders of magnitude greater than the tiling of the second matrix (in1), as the tiling of the second matrix will reuse the already compiled tiling program for the first one. By calculating the compilation time as the difference between the execution times of the first and subsequent executions, we obtain that the tiling kernel requires 296 ms to compile, regardless of the matrix dimensions. Whilst, the tiling execution time ranges from 0.351 ms for the 256x256 matrix to 2.256 ms for a 4096x4096 matrix. On the other hand, the MatMul kernel requires 620 ms for compilation, with executing times ranging from 0.328 ms to 4.783 ms for the same matrix sizes.
@@ 5.3 Performance of the MatMul kernel
3(b) shows the speedup of the MatMul kernel for different sets of cores, ranging from 1 core to 64 cores, corresponding respectively to the grid core selection (0,0)00(0,0)( 0 , 0 ) and (8,8)88(8,8)( 8 , 8 ).
@@ 5.4.1 Performance comparison
The obtained performance results have been compared against the PyTorch MatMul kernel execution on three SoA architectures: two GPUs, the NVIDIA A100 SXM4 40GB (w. peak BF16 throughput of 312 TFLOPs) and the NVIDIA V100S PCIe 32GB (w. peak FP16 throughput of 112 TFLOPs), and the Intel Sapphire Rapids server processor (w. peak BF16 throughput of 229 TFLOPs) [^2] Calculated considering AMX’s 1024 FLOPS/Cycle, 112 cores and 2.0 GHz frequency. The comparison is shown in 5(a). The Sapphire Rapids Intel processor is equipped with the Intel Advanced Matrix Extensions (AMX) to optimize MatMul execution in BF16 and INT8 formats, which are leveraged by the PyTorch’s MatMul kernel. The NVIDIA GPUs obtain remarkable performance thanks to the introduction of Tensor Cores, which accelerate the MatMul computation using BF16 and FP16 formats.